OpenStudy (anonymous):

4. Finish and balance the following chemical equations: a. H2SO4 + Al(OH)3  b. Al + Cl2  c. Ag + HCl  d. MgCl2 

OpenStudy (nechirwan):

3 H2So4 + 2 Al(OH)3 = Al2(So4)3 + 6 H2O 2 Al(s) + 3 Cl2(g) = 2 AlCl3(s) 2Ag + 2HCl -----> 2AgCl + H2 MgCl2 → Mg + Cl2